Deal of buy 1 pair boots, get 2 pair free. Person A boots $400, Person B boots $350, Person C boots $200 with sales tax of 9.5%, how much should each person pay for it to be fair?
Answer by JulietG(1812)

That depends. Are you looking for each person to contribute an equal amount? That doesn't seem quite fair since A's boots are more expensive. Is there more information?
.
In a real-life situation, boots A would be charged and B & C would be free (always the higher priced item is charged). In that case, they should split it equally, $133.33 each.
.
If, however, you're looking for them to pay based on the value of their boots, we still need to know WHICH boots are free. Again, assuming that A pays and B&C are free:
The total value of the boots are 400+350+219 (only C pays sales tax???) = 969
Take the value of each and divide by the total value to see each person's share.
A: 400/969 = 41.3% (rounded)
B: 350/969 = 36.1%
C: 219/969 = 22.6%